# Constraint Learning for Parametric Point Cloud

### v1

@@ 4.3 Classification
The classification results on the Parametric 20000 dataset are presented in Table 1, where our CstNet outperformed other methods across both metrics. Additionally, the instance accuracy of all methods is below 90 %, with a notable performance gap between them, indicating that this dataset poses challenges and effectively differentiates between methods.
@@
We conducted evaluations on the CAD shapes dataset MCB [16], with results presented in Table 2. Our CstNet achieved the highest scores across all metrics on MCB-A, but performed less effectively on MCB-B. This discrepancy may stem from the composition of MCB: MCB-A includes shapes from TraceParts (69.27 %), 3D Warehouse (21.70 %), and GrabCAD (9.03 %), while MCB-B is limited to shapes from 3D Warehouse and GrabCAD. TraceParts provides data from manufacturing companies, whereas 3D Warehouse and GrabCAD contain shapes for general usages [16]. As the CST-PCD module in CstNet is pre-trained on ABC, which primarily comprises mechanical components, this pre-training may lead to poor constraint prediction performance on MCB-B, thereby influencing the classification results.
@@ 4.4 Segmentation
We conducted experiments on the mechanical segmentation dataset 360 Gallery [19], which contains point clouds with segmentation labels based on CAD modeling operations. However, this labeling approach presents disadvantage for point cloud segmentation. For example, a cube in this dataset has four faces labeled as "ExtrudeSide" and two faces labeled as "ExtrudeEnd", which poses an unrealistic task for point cloud networks. Additionally, the distribution of points across different labels is highly imbalanced. For instance, points labeled as "ExtrudeSide" accounts for 45.94 %, while those labeled as "RevolveEnd" make up only 0.12 %. Despite these limitations, to our knowledge, 360 Gallery is currently the only publicly available point cloud dataset for CAD shapes segmentation task.
The segmentation results are presented in Table 3, where our method outperforms the state-of-the-art methods on most metrics. The results also highlight the limitations of this 360 Gallery in terms of point cloud segmentation. Apart from "ExtrudeSide" and "ExtrudeEnd", none of the models were able to differentiate the other segmentation categories except DGCNN managed to predict "RevolveSide" with a low mIoU.
Table 3: Part Segmentation Results on 360 Gallery. PN: PointNet. PN2: PointNet++. DGN: DGCNN. 3DN: 3DGCN. OA: overall accuracy %. mIoU: instance mIoU. mIoU*: class mIoU. ESide: extrude side mIoU, EEnd: extrude end mIoU. Other segmentation categories’ mIoU are all 0 except RevolveSide mIoU of DGCNN is 9.60, more datails in supplementary.
@@
A portion of the segmentation results on 360 Gallery test set are shown in Table 4, where our CstNet shows segmentation results closest to the Ground Truth. PointNet’s segmentation differs from other methods, for instance, in the third and fourth rows, PointNet mistakenly segments a single primitive into two different categories. This discrepancy likely arises from the model’s inability to effectively capture local features.
@@ 4.5 Ablation Studies
For each component of constraint, we evaluated their importance by disable one at a time. This operation is conducted by setting it as a zero tensor before feeding into the model. The results are presented in Table 5, which show that all three components contribute to improve the performance of our CstNet.
To assess the effectiveness of SurfaceKNN, we compared it with KNN. From Table 5, it can be concluded that in most cases, our SurfaceKNN outperforms KNN. However, when Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
@@ 5 Conclusions
In this paper, we studied the CAD shapes and introduced the constraints expression which suitable for deep learning methods. Afterward, we proposed the CstNet for extracting and learning constraint features. Additionally, we built up a multimodal CAD shape classification dataset Parametric 20000. Experimental results on both public datasets and proposed dataset demonstrate that our CstNet achieves state-of-the-art performance across multiple metrics. Comprehensive ablation studies further validate that each component of constraint and SurfaceKNN contributes to improve the performance of CstNet.
CAD shapes possess numerous attributes. In this paper, we analyze the constraints from functional perspective. In future work, we plan to investigate additional aspects such as part design and manufacturing processes, to achieve a deeper understanding of CAD shapes.

### v2

@@ 4.3 Classification
The classification results on the Parametric 20000 dataset are presented in Tab. 1, where our CstNet outperformed other methods across both metrics. Additionally, the instance accuracy of all methods is below 90 %, with a notable performance gap between them, indicating that this dataset poses challenges and effectively differentiates between methods.
@@
We conducted evaluations on the CAD shape dataset MCB [16], with results presented in Tab. 2. Our CstNet achieved the highest scores across all metrics on MCB-A, but performed less effectively on MCB-B. This discrepancy may stem from the composition of MCB: MCB-A includes shapes from TraceParts (69.27 %), 3D Warehouse (21.70 %), and GrabCAD (9.03 %), while MCB-B is limited to shapes from 3D Warehouse and GrabCAD. TraceParts provides data from manufacturing companies, whereas 3D Warehouse and GrabCAD contain shapes for general usage [16]. As the CST-PCD module in CstNet is pre-trained on ABC, which primarily comprises mechanical components, this pre-training may lead to poor constraint prediction performance on MCB-B, thereby influencing the classification results.
@@ 4.4 Segmentation
We conducted experiments on the mechanical segmentation dataset 360 Gallery [19], which contains point clouds with segmentation labels based on CAD modeling operations. However, this labeling approach presents disadvantages for point cloud segmentation. For example, a cube in this dataset has four faces labeled as "ExtrudeSide" and two faces labeled as "ExtrudeEnd", which poses an unrealistic task for point cloud networks. Additionally, the distribution of points across different labels is highly imbalanced. For instance, points labeled as "ExtrudeSide" account for 45.94 %, while those labeled as "RevolveEnd" make up only 0.12 %. Despite these limitations, to our knowledge, 360 Gallery is currently the only publicly available point cloud dataset for CAD shape segmentation tasks.
The segmentation results are presented in Tab. 3, where our method outperforms the state-of-the-art methods on most metrics. The results also highlight the limitations of 360 Gallery in terms of point cloud segmentation. Apart from "ExtrudeSide" and "ExtrudeEnd", none of the models were able to differentiate the other segmentation categories except DGCNN managed to predict "RevolveSide" with a low mIoU.
Table 3: Part Segmentation Results on 360 Gallery. PN: PointNet. PN2: PointNet++. DGN: DGCNN. 3DN: 3DGCN. OA: overall accuracy %. mIoU: instance mIoU %. mIoU*: class mIoU %. ESide: Extrude Side mIoU %, EEnd: Extrude End mIoU %. Other segmentation categories’ mIoU are all 0 except RevolveSide mIoU of DGCNN is 9.60 %, more datails in supplementary.
@@
A portion of the segmentation results on 360 Gallery test set are shown in Tab. 4, where our CstNet shows segmentation results closest to the Ground Truth. PointNet’s segmentation differs from other methods, for instance, in the third and fourth rows, PointNet mistakenly segments a single primitive into two different categories. This discrepancy likely arises from the model’s inability to effectively capture local features.
@@ 4.5 Ablation Studies
For each component of constraint, we evaluated their importance by disabling one at a time. This operation is performed by setting it as a zero tensor before feeding into the model. The results are presented in Tab. 5, which show that all three components contribute to improving the performance of our CstNet.
To assess the effectiveness of SurfaceKNN, we compared it with KNN. From Tab. 5, it can be concluded that in most cases, our SurfaceKNN outperforms KNN. However, when the Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
@@ 5 Conclusions
In this paper, we studied the CAD shapes and introduced the constraints expression which is suitable for deep learning methods. Afterward, we proposed the CstNet for extracting and learning constraint features. Additionally, we built up a multimodal CAD shape classification dataset Parametric 20000. Experimental results on both the public datasets and proposed dataset demonstrate that our CstNet achieves state-of-the-art performance across multiple metrics. Comprehensive ablation studies further validate that each component of constraint and SurfaceKNN contributes to improve the performance of CstNet.
CAD shapes possess numerous attributes. In this paper, we analyze the constraints from a functional perspective. In future work, we plan to investigate additional aspects such as part design and manufacturing processes, to achieve a deeper understanding of CAD shapes.